Carla invested $28,000, part at 14% and part at 9%. If the total interest at the end of the year is $3,370, how much did she invest at 14%?
Answer by checkley71(8403)

.14x+.09(28000-x)=3370 or .14x+2520-.09x=3370 or .05x=850 or x=850/.05 or
x=17000 invested @ 14% & 28000-17000=11000 @ 9%